OpenStudy (anonymous):

segments AB, BC and AC are tangents to a circle. If AB=15, BC=16, AC= 13 Find AE

OpenStudy (anonymous):

what is E

OpenStudy (anonymous):

theres a diagram in the attachment!! its like...the point of tangency of AC

OpenStudy (anonymous):

but how do you know that ABE is a right triangle?

OpenStudy (anonymous):

@rsvitale, how do you know ABE is a right triangle?

OpenStudy (anonymous):

lol...

OpenStudy (anonymous):

Okay quick question not involving math... Is your name Ivory? Mines ivy

OpenStudy (anonymous):

XD yes...my name is Ivory...=D

OpenStudy (anonymous):

OMG!! Thats effing awesome!! LOL ^_^ I always have people call me Ivory instead of Ivy :D

OpenStudy (anonymous):

BUT PEOPLE ALWAYS CALL ME IVY!! XD anyways...this is the answer to the question...but i still dont understand it... "the perpendicular bisector of a chord contains the centre of the circle, a tangent to a circle is perpendicular to the radius at the point of tangency and the tangent segments to a circle, from any external point, are congruent."

OpenStudy (anonymous):

RLY!?!?!!?! Thats fricking awesome =3 Never met anyone with anything close to my name ;D ;D

OpenStudy (anonymous):

point of tangency of AB=F, point of tangency of BC=G AE=AF=a FB=BG=c GC=EC=b

OpenStudy (anonymous):

a+c=15 a+b=13 c+b=16

OpenStudy (anonymous):

you can solve that for a=AE

OpenStudy (anonymous):

OOOOOOOOOOOOOOOOOOOOOOOOOOOOOO!!!!!!!!!! omgomgomg!!! that is genius!! thank you!!!!!

OpenStudy (anonymous):

youre welcome, sorry for the initial mistake :)

OpenStudy (anonymous):

thats ok!!! thank you soo much!

OpenStudy (anonymous):

@rsvitale, well done

OpenStudy (anonymous):

thanks